OpenStudy (anonymous):

differential equation using separable variable method of dy/y=dx/x^2(1-x^2)=0

OpenStudy (anonymous):

\[\frac{dy}{y}=\frac{dx}{x^2(1-x^2)}\] Integrate both sides: \[\int\frac{dy}{y}=\int\frac{dx}{x^2(1-x^2)}\] Split the right hand side into partial fractions: \[\frac{1}{x^2(1-x^2)}=\frac{A}{x}+\frac{B}{x^2}+\frac{C}{1-x}+\frac{D}{1+x}\\ 1=Ax(1-x^2)+B(1-x^2)+Cx^2(1+x)+Dx^2(1-x)\\ 1=Ax-Ax^3+B-Bx^2+Cx^2+Cx^3+Dx^2-Dx^3\\ 1=(-A+C-D)x^3+(-B+C+D)x^2+Ax+B\] yielding the system \[\begin{cases} -A+C-D=0\\-B+C+D=0\\A=0\\B=1 \end{cases}~\Rightarrow~A=0,B=1,C=\frac{1}{2},D=\frac{1}{2}\] So your equation becomes \[\int\frac{dy}{y}=\int\left(\frac{1}{x^2}+\frac{1}{2(1-x)}+\frac{1}{2(1+x)}\right)~dx\\ \int\frac{dy}{y}=\int\frac{dx}{x^2}+\frac{1}{2}\int\frac{dx}{1-x}+\frac{1}{2}\int\frac{dx}{1+x}\\ \ln|y|=-\frac{1}{x} +\frac{1}{2}\ln|1-x|+\frac{1}{2}\ln|1+x|+C \]